OpenStudy (anonymous):

A crow wishing to quench his thirst came to a vessel, which contained 37 cubic inches of water. The crow being unable to reach the water, picked up several small stones. Each stone was three quarters of a cubic inch in size. Then, he dropped the stones into the vessel until the water came to the top of it. If the vessel can hold 97 cubic inches of water, the number of stones, which the crow dropped into the vessel, was

OpenStudy (campbell_st):

well find the difference between the current volume and the maximum volume difference = max - current let x = number of stones \[\frac{3x}{4} = difference\] just solve for x

OpenStudy (anonymous):

diff=97-37=60 3x/4=60 so x=80

OpenStudy (campbell_st):

thats correct